OpenStudy (anonymous):

how do i find my y coordinate? find the intersection point, x+y=2, 2x-y=1

OpenStudy (anonymous):

so far i've gotten to x=1

OpenStudy (anonymous):

should i solve for x then place that in an x spot in the other equation?

OpenStudy (anonymous):

y=1 x=1

OpenStudy (anonymous):

Yes ma'am.

myininaya (myininaya):

2x-y=1 -2(x+y=2) ---------- 0 -3y=-3 -3y=-3 y=1 of y=1 then 2x-1=1 => 2x=2=>x=1

OpenStudy (mimi_x3):

2x + 2y = 4 2x - y = 1 3y = 3 y = 1 2x - 1 = 1 2x = 2 x = 1

OpenStudy (mimi_x3):

Therefore I (1,1)